if the arithmetic mean and geometric mean of three numbers a,b and c are 19 and 15 respectively. find value of a and b when c=27.
1
Expert's answer
2021-09-30T08:07:28-0400

If the mean of "a+b+c = 19" , the sum of three numbers is "19*3 = 57."

Given that the geometric mean of the three numbers is the cube root of "(a*b*c) = 15, 15^3=3375."

Factoring 3375 yields "3*3*3*5*5*5."

If "c=27 (3*3*3)" then "a=5, b=25 \space or \space a=25, b=5."

Mean of "a+b+c = \frac{(5+25+27)}{3} = 19."

Geometric mean ="\frac{(5*25*27)}{3} = 15."